# LLM-ABBA: Understanding time series via symbolic approximation

### v4

@@ IV-D Time Series Regression Tasks
Number RMSE RMSE RMSE RMSE RMSE RMSE RMSE RMSE RMSE RMSE
AppliancesEnergy 778 1.73 2.09 1.74 2.43 2.43 2.43 2.34 2.02 2.11 2.29
HouseholdPowerConsumption1 1717 377.02 377.20 377.20 398.01 398.05 398.05 228.83 228.78 228.67 132.80
HouseholdPowerConsumption2 1717 27.64 27.71 27.73 36.63 36.71 36.69 24.54 24.56 24.51 32.61
BenzeneConcentration 3037 4.01 4.00 4.00 5.57 5.56 5.56 4.03 4.03 4.03 0.64
BeijingPM10Quality 970 66.16 66.07 66.07 93.25 93.26 93.26 65.25 65.25 65.24 93.14
BeijingPM25Quality 970 54.16 54.16 54.16 76.75 76.73 76.73 53.50 53.49 53.49 59.50
LiveFuelMoistureContent 5689 20.56 20.56 20.56 29.32 29.33 29.32 20.94 20.88 20.85 29.41
FloodModeling1 969 0.00 0.00 0.00 0.05 0.05 0.05 0.37 0.36 0.36 0.00
FloodModeling2 979 0.00 0.00 0.00 0.05 0.04 0.04 0.40 0.39 0.39 0.01
FloodModeling3 948 0.00 0.00 0.00 0.06 0.05 0.05 0.41 0.37 0.39 0.00
AustraliaRainfall 4740 4.36 4.36 4.36 6.05 6.01 6.02 4.31 4.28 4.30 8.12
PPGDalia 12298 9.32 9.32 9.32 12.54 12.50 12.52 9.04 9.02 9.03 9.92
IEEEPPG 8971 17.06 17.00 17.04 22.59 22.53 22.55 17.15 17.12 17.16 23.90
BIDMC32HR 9423 6.73 6.98 6.71 12.02 11.98 12.04 8.24 8.21 8.23 9.42
BIDMC32RR 9412 1.77 1.74 1.76 2.64 2.61 2.62 2.09 2.06 2.08 3.02
BIDMC32SpO2 5537 2.90 2.85 2.89 3.82 3.79 3.81 2.95 2.91 2.93 4.45
NewsHeadlineSentiment 5537 0.07 0.07 0.07 0.13 0.13 0.13 0.11 0.11 0.11 0.14
NewsTitleSentiment 5537 0.07 0.07 0.07 0.13 0.13 0.13 0.11 0.11 0.11 0.14
Covid3Month 227 0.02 0.02 0.02 0.11 0.11 0.11 0.45 0.44 0.44 0.04
@@
Experimenting on the TSER benchmark archive [2], the empirical results are shown in Table VII, in which for 15 out of 19 use-cases, LLM-ABBA outperforms the machine learning SOTA results. We believe that LLM-ABBA can exploit the semantic information hiding beneath the time series in the task of time series regression. ABBA is able to provide COPs to LLMs by compressing and digitizing time series to symbols, which finally results in the change of embedding space by using adaption fine-tuning methods.
IV-E Time Series Forecasting Tasks
@@
Figure 7: Visualization of input-168-predict-24 results on ETTh1 using LLM-ABBA.
TABLE VIII: Full comparison of results for the prediction task on 4 time series prediction datasets.
Data Predictor Symbols Llama2-7B Mistreal-7B
Informer
[2]
Time-LLM
[6]
TimeMixer
[7]
@@
For time series forecasting, we experimented on 4 well-established benchmarks: ETT datasets (including 4 subsets: ETTh1, ETTh2, ETTm1, ETTm2) [45, 46]. Details of the implementation and datasets can be found in Table III. The input length of the time series is $168$, and we use three different prediction horizons $H\in\{24,96,168\}$. The evaluation metrics include MSE and MAE.
@@
IV-F QLoRA Fine-Tuning
Because the low rank of adapter fine-tuning will influence the efficiency of passing information [17, 30] from the previous layer, we use different low rank settings of QLoRA on the corresponding tasks during the fine-tuning progress. But for time series regression and prediction tasks, we select $r\in\{16,46,256\}$ for the corresponding data input. We find that there is no obvious over-fitting problem, and more tunable parameters are not able to improve the performance of LLM-ABBA.
In medical time series domains, ptb-db and MIT-BIH arrhythmia data sets are mostly used. EEG eye state data set has two categories,and because of its high complexity, the accuracy always stays at around $60\%$. EEG eye state data and MIT-BIH has more than one channel, which indicates that LLM-ABBA might have the ability to process complicate features across channels. Table VI presents the full medical time series classification results using LLM-ABBA.
LLM-ABBA achieves comparable time series prediction results to the SOTAs, and there is no over-fitting in these tasks when using different low rank $r$. Because ABBA tends to symbolize trends and altitudes of the time series signals, LLM-ABBA always strengthens the vibration of predicted time series segments which can be seen in Figure 7.
IV-G Semantic consistency
When using pretrained tokens as the input symbols, fine-tuning on no language content (such as time series signals) will generally bring semantic loss to LLMs. Therefore, we use ASCII codes to generate new symbols by adding more digits and expanding the used alphabet table. Following the same fine-tuning process to the above experiment settings, we compute the forecasting performance by fine-tuning on Mistral-7B. Compared to Table VIII, Table IX shows that the difference is not noticeable.
@@ IV-G Semantic consistency
Data Predictor Symbols Mistreal-7B
@@
V Limitations
@@
ABBA is assessed carefully via performance profiles with respect to its reconstruction evaluated via 2-norm, DTW, and their respective differenced measures, which shows competitive performance against the SOTA STSA methods (e.g., SAX). Previous STSA methods have been applied in various data mining applications, e.g., EEG signal analysis [47] and Internet of Things [48]. ABBA also shows improved performance in the anomaly detection method TARZAN by simply replacing SAX methods [13, 14].
@@
LLMs can understand generated symbols of ABBA. Each data sample can be approached by symbols, and each used symbol has a specific meaning which presents one node of the internal COP of time series data. LLM-ABBA performs well not only on time series classification tasks, but also on time series regression tasks (as seen in Table V and Table VII). Because these symbolic series have a logic chain that can represent the trend of time series samples, LLMs are able to learn the trend of time series via adapter fine-tuning methods. As shown in Figure 7, by using the inverse symbolization process of ABBA, LLMs can predict the trend of time series signals, and these predicted parts have a smaller drift. Therefore, time series forecasting tasks also can demonstrate these findings. ABBA perfectly approximates time series via symbolic series, but because LLMs are born with hallucination, more generated contents would contain more “hallucination knowledge”. Therefore, LLM-ABBA performs better on short-term time series prediction tasks and time series regression tasks.
Our proposed FAPCA strategy for ABBA still cannot fully guarantee a complete removal of cumulative error arising from the previous mistaken symbols from the recovery, because an incorrect $\text{{len}}_{i}$ might leads to a minor shift if certain symbols are replaced improperly. Additionally, as a congenital defect of LLMs, hallucination cannot be addressed in this work, and the vibration or adverse response of predicted sequences can still have negative effects on final performance. Moreover, after using ABBA to transform time series, most LLMs only can support up to 4,096 tokens, which fundamentally prohibits long-term time series analysis tasks.
VI Conclusion
@@
In this paper, we propose LLM-ABBA for time series classification, regression, and forecasting tasks. We discuss how to seamlessly integrate time series symbolization with LLMs and enhance its performance. In theory, we analyze the reconstruction error of ABBA symbolization, how it relates to the dominant parameters, and the congenital defect of LLM-ABBA. To mitigate the drift phenomenon of time series, we introduce the FAPCA method to improve ABBA symbolization. The empirical results demonstrate our method achieves a performance comparable to the SOTA on classification and regression tasks. In terms of convenience and universality, LLM-ABBA improves the multi-modality of LLMs on time series analysis. We believe the potential of ABBA extends to other time series applications, which will be left as future work.

### v5

@@ IV-E Time series regression tasks
Number RMSE RMSE RMSE RMSE RMSE RMSE RMSE RMSE RMSE RMSE RMSE
AppliancesEnergy 778 1.73 2.09 1.74 2.43 2.43 2.43 2.34 2.02 2.11 2.81 2.27
HouseholdPowerConsumption1 $17$ 377.02 377.20 377.20 398.01 398.05 398.05 228.83 228.78 228.67 188.59 146.96
HouseholdPowerConsumption2 $17$ 27.64 27.71 27.73 36.63 36.71 36.69 24.54 24.56 24.51 35.99 35.61
BenzeneConcentration 3037 4.01 4.00 4.00 5.57 5.56 5.56 4.03 4.03 4.03 5.63 3.57
BeijingPM10Quality 970 66.16 66.07 66.07 93.25 93.26 93.26 65.25 65.25 65.24 96.49 93.05
BeijingPM25Quality 970 54.16 54.16 54.16 76.75 76.73 76.73 53.50 53.49 53.49 64.01 60.44
LiveFuelMoistureContent 5689 20.56 20.56 20.56 29.32 29.33 29.32 20.94 20.88 20.85 34.92 33.69
FloodModeling1 969 0.002 0.002 0.002 0.05 0.05 0.05 0.37 0.36 0.36 0.01 0.004
FloodModeling2 979 0.002 0.003 0.002 0.05 0.04 0.04 0.40 0.39 0.39 0.01 0.01
FloodModeling3 948 0.002 0.002 0.002 0.06 0.05 0.05 0.41 0.37 0.39 0.01 0.01
AustraliaRainfall 4740 4.36 4.36 4.36 6.05 6.01 6.02 4.31 4.28 4.30
PPGDalia 12298 9.32 9.32 9.32 12.54 12.50 12.52 9.04 9.02 9.03 15.09 15.24
IEEEPPG 8971 17.06 17.00 17.04 22.59 22.53 22.55 17.15 17.12 17.16 34.92 33.69
BIDMC32HR 9423 6.73 6.98 6.71 12.02 11.98 12.04 8.24 8.21 8.23 8.01 10.62
BIDMC32RR 9412 1.77 1.74 1.76 2.64 2.61 2.62 2.09 2.06 2.08 4.30 4.04
BIDMC32SpO2 5537 2.90 2.85 2.89 3.82 3.79 3.81 2.95 2.91 2.93 4.86 5.10
NewsHeadlineSentiment 5537 0.07 0.07 0.07 0.13 0.13 0.13 0.11 0.11 0.11 0.15 0.15
NewsTitleSentiment 5537 0.07 0.07 0.07 0.13 0.13 0.13 0.11 0.11 0.11 0.14 0.14
Covid3Month 227 0.02 0.02 0.02 0.11 0.11 0.11 0.45 0.44 0.44 0.04 0.04
@@
Figure 7: The RMSE ranks for 23 Regressors on 19 TSER datasets. There are 20 of them used in [15].
Experimenting on the TSER benchmark archive [45], the empirical results are shown in TABLE VII, in which for 12 out of 19 use-cases, RoBERTalarge-ABBA outperforms the machine learning SOTA results, such as DrCIF [15] and FreshPRINCE [35]. We believe that LLM-ABBA can exploit the semantic information hiding beneath the time series in the task of time series regression. Using the same post-hoc two-tailed Nemenyi test to [45], Figure 7 shows that ReBERTalarge-ABBA is the most accurate algorithm with an average rank of $2.72$. The critical difference (CD) is $7.356$. There are 23 regressors, and 20 of them come from [15]. ABBA is able to provide COPs to LLMs by compressing and digitizing time series to symbols, which finally results in the change of embedding space by using adaption fine-tuning methods.
IV-F Time series forecasting tasks
@@
Figure 8: Visualization of input-168-predict-24 results on ETTh1 using Mistreal-7B-ABBA.
TABLE VIII: Full comparison of results for the forecasting task on 4 time series forecasting datasets.
Data Predictor Symbols Llama-2-7B-ABBA Mistreal-7B-ABBA Informer [45] Time-LLM [23] TimeMixer [49]
@@
For time series forecasting, we experimented on 4 well-established benchmarks: ETT datasets (including 4 subsets: ETTh1, ETTh2, ETTm1, ETTm2) [52]. Details of the implementation and datasets can be found in TABLE II. The input length of the time series is $168$, and we use three different forecasting horizons $H\in\{24,96,168\}$. The evaluation metrics include MSE and MAE.
@@
IV-G QLoRA fine-tuning
Because the low rank of adapter fine-tuning will influence the efficiency of passing information [10, 26] from the previous layer, we use different low rank settings of QLoRA on the corresponding tasks during the fine-tuning progress. But for time series regression and forecasting tasks, we select $r\in\{16,64,256\}$ for the corresponding data input. We find that there is no obvious over-fitting problem, and more tunable parameters are not able to improve the performance of LLM-ABBA. In medical time series domains, ptb-db and MIT-BIH arrhythmia data sets are mostly used. EEG eye state data set has two categories, and because of its high complexity, the accuracy always stays at around $60\%$. EEG eye state data and MIT-BIH has more than one channel, which indicates that LLM-ABBA might have the ability to process complicate features across channels. TABLE VI presents the full medical time series classification results using LLM-ABBA.
LLM-ABBA achieves comparable time series forecasting results to the SOTAs, and there is no over-fitting in these tasks when using different low rank $r$. Because ABBA tends to symbolize trends and altitudes of the time series signals, LLM-ABBA always strengthens the vibration of predicted time series segments which can be seen in Figure 8.
IV-H FAPCA benefits the time series forecasting
When focusing solely on the symbolic approximation in time series forecasting tasks, Figure 9 illustrates the difference between traditional ABBA and ABBA with FAPCA. If an incorrect replacement of a previous symbol occurs in the subsequent reconstruction, traditional ABBA is prone to the accumulation of errors. In contrast, FAPCA effectively mitigates this issue, particularly in long-term forecasting scenarios such as the Input-168-Predict-168 task.
(a) Traditional ABBA.
(b) ABBA with FAPCA.
Figure 9: Visualization of Input-168-Predict-168 with the use of the traditional ABBA and the ABBA with FAPCA on ETTh1 data.
IV-I Semantic consistency
When using pretrained tokens as the input symbols, fine-tuning on non-linguistic data (e.g., time series signals) may cause semantic drift in LLMs. To mitigate this issue, we construct a new set of symbols to extend ASCII codes by adding more digits and expanding the alphabet table to be used. Following the same fine-tuning configuration as in the previous experiments, we assess the forecasting performance by fine-tuning Mistral-7B. As shown in TABLE IX, the results are largely consistent with those in TABLE VIII, indicating that the semantic loss is negligible.
@@ IV-I Semantic consistency
Data Predictor Symbols Mistreal-7B-ABBA
@@
IV-J Trade-off between runtime and performance
Figure 10 shows the average rank RMSE against the run time (on a log scale) for the nine regressors: RoBERTalarge-ABBA, Mistreal-7B-ABBA, Llama-2-7b-ABBA, ROCKET [9], MultiROCKET [46], InceptionE [19], DrCIF [15], FreshPRINCE [35], and CNN [18]. We see a direct trade-off between runtime and performance. All algorithms run on a single thread CPU except for RoBERTalarge-ABBA, Mistreal-7B-ABBA, and Llama-2-7b-ABBA, which ran on a GPU. The total run time of ReBERTalarge-ABBA on 19 TSER datasets is shorter than that of InceptionE [19], DrCIF [15], and FreshPRINCE [35]. But because of the huge weight of LLMs, the learning time and the inference of LLM-ABBA are inevitably longer than that of other traditional machine learning methods in [45, 15].
Figure 10: Run time in seconds (log scale over 19 TSER datasets) plotted against average rank for RMSE.
V Discussion
@@
ABBA is assessed carefully via performance profiles with respect to its reconstruction evaluated via 2-norm, DTW, and their respective differenced measures, which show competitive performance against the SOTA STSA methods (e.g., SAX). Previous STSA methods have been applied in various data mining applications, e.g., EEG signal analysis [20] and Internet of Things [21]. ABBA also shows improved performance in the anomaly detection method TARZAN by simply replacing SAX methods [12, 3].
@@
LLMs can understand the generated symbols of ABBA. Each data sample can be approached by symbols, and each used symbol has a specific meaning that presents one node of the internal COPs of time series data. LLM-ABBA performs well not only on time series classification tasks, but also on time series regression tasks (as seen in TABLE V and TABLE VII). Because these symbolic series have a logic chain that can represent the trend of time series samples, LLMs are able to learn the trend of time series via adapter fine-tuning methods. As shown in Figure 8, by using the inverse symbolization process of ABBA, LLMs can predict the trend of time series signals, and these predicted parts have a smaller drift. Therefore, time series forecasting tasks also can demonstrate these findings. ABBA perfectly approximates time series via symbolic series, but because LLMs are born with hallucination, more generated contents would contain more “hallucination knowledge”. Therefore, LLM-ABBA performs better on short-term time series forecasting tasks and time series regression tasks.
VI Conclusion and Limitations
@@
In this paper, we propose LLM-ABBA for time series classification, regression, and forecasting tasks. We discuss how to seamlessly integrate time series symbolization with LLMs and enhance its performance. In theory, we analyze the reconstruction error of ABBA symbolization, how it relates to the dominant parameters, and the congenital defect of LLM-ABBA. To mitigate the drift phenomenon of time series, we introduce the FAPCA method to improve ABBA symbolization. The empirical results demonstrate our method achieves a performance comparable to the SOTA on classification and regression tasks. In terms of convenience and universality, LLM-ABBA improves the multi-modality of LLMs on time series analysis.
The proposed FAPCA strategy for ABBA cannot fully guarantee a complete removal of cumulative error arising from the previous mistaken symbols from the recovery. Additionally, the hallucination of LLMs cannot be addressed in this work, and the vibration or adverse response of predicted sequences can still have negative effects on final performance. Moreover, most LLMs only can support up to 4,096 tokens, which fundamentally prohibits long-term time series analysis tasks. Lastly, the learning time and the inference speed of LLM-ABBA are inevitably slower than most traditional machine learning methods.